# Explanation Bias is a Product: Revealing the Hidden Lexical and Position Preferences in Post-Hoc Feature Attribution

### v1

@@ Abstract
Good quality explanations strengthen the understanding of language models and data. Feature attribution methods, such as Integrated Gradient, are a type of post-hoc explainer that can provide token-level insights. However, explanations on the same input may vary greatly due to underlying biases of different methods. Users may be aware of this issue and mistrust their utility, while unaware users may trust them inadequately. In this work, we delve beyond the superficial inconsistencies between attribution methods, structuring their biases through a model- and method-agnostic framework of three evaluation metrics. We systematically assess both the lexical and position bias (what and where in the input) for two transformers; first, in a controlled, pseudo-random classification task on artificial data; then, in a semi-controlled causal relation detection task on natural data. We find that lexical and position biases are structurally unbalanced in our model comparison, with models that score high on one type score low on the other. We also find signs that methods producing anomalous explanations are more likely to be biased themselves.
@@ 1 Introduction
Figure 1: Bias distributions of top-1 token explanations from ModernBERT classifiers over positions (left; 0–19) and lexical elements (right; "," and ".") in the input. Here we compare two feature attribution methods: Vanilla Gradient (top); Integrated Gradient (bottom). The dotted line indicates the position where we would expect the predictions to be in the absence of bias.
@@
Feature attribution is a post-hoc explanation technique that aims to determine the input tokens that contributed the most to a language model’s prediction. Different attribution methods have been proposed, but they tend to exhibit inconsistent token explanations even on the same input and given the same backbone model (neely2022song). A major problem that comes with attribution inconsistency is the lack of adequate domain knowledge: methods are either adopted without considering the risks of unfaithful explanations (i.e. that do not reflect the true decision mechanisms of the model), or are coarsely deemed unreliable and discarded altogether. Their inconsistent explanation behaviour can manifest itself through various symptoms that reflect some underlying systematic deviation, i.e. a bias, that resides in or surfaces through the attribution method. For example, lyu2024towards suggests that, when methods attribute importance to words such as determiners (the) over adjectives (great), the model is either relying on function words, or the method is not faithfully reflecting the model’s reliance on content words. Since this type of bias concerns lexical information, we refer to it as lexical bias.
@@
However, lexical information might only be a part of the full story, as positional information likely plays a role in bias, too. Figure 1 illustrates the bias distributions for token position and token type of two explanation methods on the same classification task on artificial sequences of periods and commas. Vanilla Gradient (simonyan2014deep) is very biased towards the start of the input texts (index 0). In contrast, the more elaborate Integrated Gradient (sundararajan2017axiomatic) seems more balanced towards position but shows some lexical bias by preferring "." over ",".
@@
In this paper, we study the sources of bias. We argue that lexical and position biases are a product of the attribution method, the model type and the random seed. In this light, we propose an evaluation framework to investigate the biases exhibited by six different attribution methods combined with a traditional and a recent BERT-based classifier. We assess how these transformers, which differ in pre-training, architectural properties and size, suffer from bias: is the modern BERT version less affected? And to what extent are position and lexical bias related? By systematically structuring bias symptoms, we aim to better understand the underlying causes of method–method inconsistencies (also known as disagreement) to help users of attribution methods make more informed decisions. We first experiment in a controlled setting with artificial data, followed by a semi-controlled setting on English causal relation detection.
@@ 2 Related Work
Attribution methods are inconsistent compared to one another (krishna2022disagreement; neely2022song; kamp-etal-2023-dynamic), which makes it challenging to interpret a backbone model. Possible underlying biases, such as the preference for certain lexical classes and their position in the input, may not only affect model understanding but also the cognitive processing of the presented explanations. For example, jacovi-etal-2023-neighboring carry out a user study showing that lexical and positional information of highlighted tokens are indicative of the way humans value explanations as important.
@@
Lexical bias in attribution methods has been explored in a handful of tasks, such as sentiment analysis and deception detection (lai2019many), and reading comprehension-based question answering (ramnath2020towards). kamp-etal-2024-role investigate the disagreement among attribution methods by mapping explanations to lexical classes. They find, on a natural language inference task, that methods may have different biases towards e.g. nouns, determiners and punctuation but still converge to the same syntactic phrase. They speculate that position bias might be induced by the proximity to an important token. The problem is related to the work of ko-etal-2020-look; ben2024impact, who show that the position of signal words in the input induces a bias towards a specific target class. xiao2023efficient find strong attention scores (sinks), on initial tokens LLM inputs, irrespective of their semantics. eshuijs-etal-2025-short instead find that in a sentiment analysis task, the first most informative word in a signal span receives the most attention in decoder-only models, hence overruling the weight of the word’s position in the span.
@@
The fact that biases are observed on specific tasks raises the question to what extent biases are task- and data-dependent; some contrasting findings by attanasio2022benchmarking versus atanasova2020diagnostic seemed to indicate a task-dependent component in the source of disagreement. We differ from previous work in that we assess both the lexical and position bias in a controlled setting to isolate the findings at best from task-related aspects. To our knowledge, we are the first to do so.
@@
A factor that may directly affect feature attribution is the complexity or size of the model. zhou-etal-2024-explaining explore the effect of training data size, showing that more data can lead to greater plausibility (i.e. alignment with human rationales) but not to greater faithfulness (i.e. alignment with model internals). Notably, they also find that a larger model (Vicuna, 13B (vicuna2023) vs. RoBERTa-large, 355M (liu2019roberta)) does not necessarily produce more plausible or faithful explanations. Conversely, heyen2024the compare LIME (ribeiro-etal-2016-trust) on four versions of DeBERTaV3 (hedebertav3) (22–304M parameters), and find that explanations do not become more plausible in larger models, but do become more faithful. The caveat is that metrics to assess faithfulness can be unreliable, too. in the same study, heyen2024the find that comprehensiveness, which approximates faithfulness as the prediction drop after removing explanation tokens from the input, is not in line with other metrics capturing greater faithfulness for the larger model; this suggests a shortcoming of this metric in specific cases.
@@ 3 Experiments on Artificial Data
We study attribution bias as the frequency in which token explanations are located at certain indices of the input (position bias) and belong to a certain word type (lexical bias). We craft the tasks in a controlled setting in such a way that this information should not play a role in solving the task (which should not be solvable). If attributions are biased, they should manifest this bias consistently across runs and datasets. In §4, we complement this randomised setting with a natural language task.
@@ 3.1 Data
The chosen lexical classes follow kamp-etal-2024-role, whose initial insights show preferences for specific parts-of-speech. We include a frequent noun (’table’) and determiner (’the’), and a period (’.’). Each token can occur 0+ times. Token positions in the sentences are allocated randomly. Class labels are random and equally distributed.
@@
Each instance is a random sequence of ’.’ and ’,’ tokens, each of which can occur 0+ times. The absence of both content and function words may highlight the lexical bias in certain attribution methods that favour punctuation. Class labels are random and equally distributed.
@@
Each instance is a unique, random ordering of the set of unique punctuation marks {".", ",", ";", ":", "!", "?", "-", "_", "(", ")", "[", "]", "", "", "/", "*", "#", "’", ’"’, "‘"}. The uniqueness of these lexical elements should facilitate faithfulness assessment (introduced in §3.4) by enabling us to measure the weight of a specific token in the input. Class labels are random and equally distributed.
@@ 3.2 Models
We compare two transformer models with a sequence classification head, namely BERT: bert-base-uncased, 110M (devlin-etal-2019-bert) and ModernBERT: ModernBERT-base, 149M (modernbert). We question whether a more recent BERT architecture suffers less from bias. After introducing the bias metrics in §3.4, we speculate further on how model properties may relate to bias.
@@ 3.3 Explanations
We compute token attribution scores for each of the input sequences. We ignore start-of-sequence and end-of-sequence tokens, such as [CLS] and [SEP], after computing the attributions in order to focus the analysis on the original 20 tokens per input. We test six different attribution methods: Partition SHAP, or PartSHAP (lundberg2017unified) LIME (ribeiro-etal-2016-trust); Vanilla Gradient, or VanGrad (simonyan2014deep), and Integrated Gradient, or IntGrad (sundararajan2017axiomatic); the latter two multiplied by the input – Grad × I and IntGrad × I shrikumar2017learning. We select the top-$k$ attributions per instance and set $k=1$, i.e. the single strongest attribution preference. This choice ensures a fair comparison with the real world task in §4, where we expect a single word signal to be indicative of the class label.
@@ 3.4 Bias Metrics
Given an attribution method, the instance-wise top-1 explanations computed on a test dataset form a frequency distribution over the lexical categories and position indices (Figure 1). These distributions may reflect a bias from the method itself. However, possible biases from pre-training and random weight initialisation would also surface in this way. To properly differentiate between bias types, we propose an analysis framework that includes three key metrics for bias quantification: (a) Bias-cons, the consistency of the frequency bias between the different random seeded runs; (b) Bias-agg, the bias of the aggregate distribution (of a specific attribution method) over 10 fine-tuned models trained with different random seeds; (c) Bias-attr, the bias that is specific to the attribution method, compared to other methods. In principle, the framework is applicable to any attribution method and any top-$k$ value. The proposed baseline is valid in a controlled setting and might need adjustment based on the nature of the task. We use a different baseline for causal relation detection in §4.
@@
We question to what extent biases are consistent across different random seeds. Since we are interested in the distribution of top-$k$ selections over input tokens, we measure the mean pairwise Jensen-Shannon distance (Bias-cons) between the normalised distributions yielded by different runs. For this, we consider the unique model pairs $(m^{\prime},m^{\prime\prime})$ formed from the set of different runs $\mathcal{M}$ and we average the JS distance over the number of combinations, i.e. $\binom{|\mathcal{M}|}{2}=\frac{|\mathcal{M}|(|\mathcal{M}|-1)}{2}$. JS ranges from 0 (least divergent/biased) to 1 (most divergent/biased). The greater the distance, the greater the impact of random initialisation on the explanations. Formally, for each model type $m\in\{\text{{BERT, ModernBERT}}\}$ and each attribution method $a\in\mathcal{A}$, let $P^{(m^{\prime})}_{m,a}$ or $P^{(m^{\prime\prime})}_{m,a}$ denote the bias distribution (over positions or lexical elements in the input) for a single run $m^{\prime}$ or $m^{\prime\prime}$ (i.e., a specific random seed):
@@
We question the extent to which attribution methods are biased in the light of a model-to-model comparison. For each model type $m$, we compute the Jensen-Shannon distance (JS) between each method $a$’s aggregate bias distribution $P_{m,a}$ and a baseline. $P_{m,a}$ is computed by summing token-level bias distributions across all runs $\mathcal{M}$ and normalising to form a probability distribution. The baseline consists of a normalised uniform distribution $U$ over the positions and lexical elements for the specific dataset, and implies the absence of bias. We refer to this distance metric as Bias-agg, the bias of the aggregated distributions. Aggregate also refers to the fact that this bias measures the conflation of both model bias and attribution method bias.
@@
If pre-training were the only source of bias, we would expect all attribution methods to exhibit a consistent bias for the same model, same random seed. This is not the case (attributions are different per method), indicating that at least one or more attribution methods are biased independently from model bias. We define attribution bias (Bias-attr) as the averaged pairwise JS distance between the target method $a$ and each of the other methods $a^{\prime}$ sampled from $\mathcal{A}$.
@@
It is unknown which of the attribution methods in $\mathcal{A}$ are the most biased. Based on the principle of reproducibility in science (popper2005logic), a finding that is confirmed through different means is more likely to be correct: can a majority consensus among methods be indicative of faithfulness? Are outlier explanations more likely to be wrong? We therefore question whether we observe similar patterns in Bias-attr and Bias-agg. In other words, are attribution methods that differ mostly from the rest (high Bias-attr) or are most similar to the rest (low Bias-attr) respectively the most and least biased with respect to a baseline expressing absence of bias (high and low Bias-agg)?
Faithfulness metrics as sanity check
@@
We assess faithfulness through two widely used metrics: sufficiency (suff) (lei2016rationalizing), which is obtained by removing the top-1 tokens (the rationale $r_{i}$) from the input $x_{i}$ and observing the change in probability towards the ground truth label; comprehensiveness (cmp) (yu2019rethinking), which is the change in probability obtained by removing the context ($x_{i}\setminus r_{i}$) around the top-1 tokens from $x_{i}$. Formally, following deyoung2020eraser:
@@
Most of the time, expecting a model to do better merely based on its size is not enough. To speculate more precisely how our models may relate to bias, we take into account the following three considerations. (i) ben2024impact directly relate a form of position bias to the type of positional embeddings in the model. Contrary to BERT, which learnt fixed positional embeddings, ModernBERT uses Rotary Positional Embeddings (RoPE): along with the larger model size, this may lead to a reduced position bias. (ii) Lexical bias may also be reduced by RoPE in ModernBERT, giving more weight to the relative position of a lexical element in the context of the sentence, hence affecting its meaning. (iii) The effect on lexical bias caused by the greater diversity of datasets used by ModernBERT in pre-training is not univocal: it may either mitigate or enforce it.

### v2

@@ Abstract
Good quality explanations strengthen the understanding of language models and data. Feature attribution methods, such as Integrated Gradient, are a type of post-hoc explainer that can provide token-level insights. However, explanations on the same input may vary greatly due to underlying biases of different methods. Users may be aware of this issue and mistrust their utility, while unaware users may trust them inadequately. In this work, we delve beyond the superficial inconsistencies between attribution methods, structuring their biases through a model- and method-agnostic framework of three evaluation metrics. We systematically assess both lexical and position bias (what and where in the input) for two transformers; first, in a controlled, pseudo-random classification task on artificial data; then, in a semi-controlled causal relation detection task on natural data. We find a trade-off between lexical and position biases in our model comparison, with models that score high on one type score low on the other. We also find signs that anomalous explanations are more likely to be biased.
@@ 1 Introduction
Figure 1: Bias distributions of top-1 token explanations from ModernBERT classifiers over positions (left; 0–19) and lexical elements (right; ‘,’ and ‘.’) in the input. Here we compare two feature attribution methods: Vanilla Gradient (top); Integrated Gradient (bottom). The dotted line indicates the position where we would expect the predictions to be in the absence of bias.
@@
Feature attribution is a post-hoc explanation technique that aims to determine which input tokens contributed most to a language model’s prediction. Different attribution methods have been proposed, but they tend to exhibit inconsistent token explanations on the same input and given the same backbone model (Neely et al., 2022). A major problem that comes with attribution inconsistency is the lack of adequate domain knowledge: methods might either be adopted without considering the risks of unfaithful explanations (i.e. explanations that do not reflect the true decision mechanisms of the model), or might be coarsely deemed unreliable and discarded altogether. Their inconsistent explanation behaviour can manifest itself through various symptoms that reflect some underlying systematic deviation, i.e. a bias, that resides in or surfaces through the attribution method. For example, Lyu et al. (2024) argue that, in cases where content words (e.g. great) are expected to be more informative than function words (e.g. the), methods that assign higher importance to function words raise the question of whether the model is truly relying on them or whether the method fails to faithfully reflect the model’s reliance on content words. We talk about a lexical bias if the attribution is systematically assigned to specific lexical elements.
@@
However, lexical information might only be a part of the full story, as positional information likely plays a role in bias, too. Figure 1 illustrates the bias distributions for token position and token type of two explanation methods on the same classification task on artificial sequences of periods and commas. Vanilla Gradient (Simonyan et al., 2014) is very biased towards the start of the input texts (index 0). In contrast, the more elaborate Integrated Gradient (Sundararajan et al., 2017) seems more balanced towards position but shows some lexical bias by preferring ‘.’ over ‘,’. [^1] We do not only consider words but any element in the lexicon to be lexical elements, including punctuation.
@@
In this paper, we study the sources of bias. We argue that lexical and position biases can be a product of the attribution method, the model type, and the random seed. In this light, we propose an evaluation framework to investigate the biases exhibited by six different attribution methods combined with a traditional and a recent BERT-based classifier. We assess how these transformers, which differ in pre-training, architectural properties and size, suffer from bias. This allows us to address questions such as: is the modern BERT version less affected? And to what extent are position and lexical bias related? By systematically structuring bias symptoms, we aim to better understand the underlying causes of method–method inconsistencies (also known as disagreement) to help users of attribution methods make more informed decisions. We first experiment in a controlled setting with artificial data, followed by a semi-controlled setting on English causal relation detection.
@@ 2 Related Work
Attribution methods are inconsistent compared to one another (Krishna et al., 2022; Neely et al., 2022; Kamp et al., 2023), which makes it challenging to interpret a backbone model. Possible underlying biases, such as the preference for certain lexical classes and their position in the input, may not only affect model understanding but also the cognitive processing of the presented explanations. For example, Jacovi et al. (2023) carry out a user study showing that lexical and positional information of highlighted tokens are indicative of the way humans value explanations as important.
@@
Lexical bias in attribution methods has been explored in a handful of tasks, such as sentiment analysis and deception detection (Lai et al., 2019), and reading comprehension-based question answering (Ramnath et al., 2020). Kamp et al. (2024) investigate the disagreement among attribution methods by mapping explanations to lexical classes. They find, on a natural language inference task, that methods may have different biases towards e.g. nouns, determiners and punctuation but often still converge to the same syntactic phrase. They speculate that position bias might be induced by the proximity to an important token. The problem is related to the work of Ko et al. (2020) and Ben Amor et al. (2024), who show that the position of signal words in the input induces a model bias towards a specific target class. Xiao et al. (2023) find strong attention scores (sinks), on initial tokens LLM inputs, irrespective of their semantics. Eshuijs et al. (2025) instead find that the first most informative word in a span receives the most attention in decoder-only models, hence overruling the weight of the word’s position in the span.
@@
The fact that biases are observed on specific tasks raises the question to what extent biases are task- and data-dependent; some contrasting findings by Attanasio et al. (2022) versus Atanasova et al. (2020) seem to indicate a task-dependent component in the source of disagreement. We differ from previous work in that we assess both lexical and position bias in a controlled setting to isolate the findings from task-related aspects. To our knowledge, we are the first to do so.
@@
Two factors that may directly affect feature attribution are the size of the fine-tuning data and the size of the model (i.e. the number of parameters). Zhou et al. (2024) explore the effect of fine-tuning data size, showing that more data usually leads to greater plausibility (i.e. alignment with human rationales) but not to greater faithfulness (i.e. alignment with model internals). Notably, they do not find a clear effect of model size (Vicuna, 13B (Chiang et al., 2023) vs. RoBERTa-large, 355M (Liu et al., 2019)) on plausibility or faithfulness; instead, the results strongly depend on the dataset and the attribution method. Conversely, Heyen et al. (2024) compare LIME (Ribeiro et al., 2016) on four versions of DeBERTaV3 (He et al., 2023) (22–304M parameters), and find that explanations do not become more plausible in larger models, but do become more faithful. The caveat is that metrics that assess faithfulness can be unreliable, too. in the same study, Heyen et al. (2024) find that comprehensiveness, which approximates faithfulness as the prediction drop after removing explanation tokens from the input, is not in line with other metrics capturing greater faithfulness for the larger model; this suggests a possible metric limitation.
@@ 3 Experiments on Artificial Data
To study attribution bias, we use the frequency in which token explanations are located at certain input indices (position bias) and belong to a certain word type (lexical bias). We craft the tasks in a controlled setting in such a way that this information should not play a role in solving the task (which should not be solvable). If attributions are biased, they should manifest this bias consistently across runs and datasets. In §4, we complement this randomised setting with a natural language task.
@@ 3.1 Data
The chosen lexical classes follow Kamp et al. (2024), whose initial insights show preferences for specific parts-of-speech. We include a frequent noun (‘table’) and determiner (‘the’), and a period (‘.’). Each token can occur zero or more times. Token positions in the sentences are allocated randomly. Class labels are random and equally distributed.
@@
Each instance is a random sequence of ‘.’ and ‘,’ tokens, each of which can occur 0+ times. The absence of both content and function words may highlight the lexical bias in certain attribution methods that favour punctuation. Class labels are random and equally distributed.
@@
Each instance is a unique, random ordering of the set of unique punctuation marks {‘.’, ‘,’, ‘;’, ‘:’, ‘!’, ‘?’, ‘-’, ‘_’, ‘(’, ‘)’, ‘[’, ‘]’, ‘{’, ‘}’, ‘/’, ‘*’, ‘#’, “”’, ‘"’, “‘”}. The uniqueness of these lexical elements should facilitate faithfulness assessment (introduced in §3.4) by enabling us to measure the weight of a specific token in the input. Class labels are random and equally distributed.
@@ 3.2 Models
We compare two transformer models with a sequence classification head, namely BERT: bert-base-uncased, 110M (Devlin et al., 2019) and ModernBERT: ModernBERT-base, 149M (Warner et al., 2024). We investigate whether a more recent BERT architecture suffers less from bias. After introducing the bias metrics in §3.4, we speculate further on how model properties may relate to bias.
@@ 3.3 Explanations
We compute token attribution scores for each of the input sequences. We ignore start-of-sequence and end-of-sequence tokens, such as [CLS] and [SEP], after computing the attributions in order to focus the analysis on the original 20 tokens per input. We test six different attribution methods: Partition SHAP, or PartSHAP (Lundberg and Lee, 2017) LIME (Ribeiro et al., 2016); Vanilla Gradient, or VanGrad (Simonyan et al., 2014), and Integrated Gradient, or IntGrad (Sundararajan et al., 2017); the latter two multiplied by the input – Grad × I and IntGrad × I Shrikumar et al. (2017). We select the top-$k$ attributions per instance and set $k=1$, i.e. the single strongest attribution preference. This choice ensures a fair comparison with the real world task in §4, where we expect a single word signal to be indicative of the class label.
@@ 3.4 Bias Metrics
Given an attribution method, the instance-wise top-1 explanations computed on a test dataset form a frequency distribution over the lexical categories and position indices (Figure 1). These distributions may reflect a bias from the method itself. However, possible biases from pre-training and random weight initialisation would also surface in this way. To properly differentiate between bias types, we propose an analysis framework that includes three key metrics for bias quantification: (a) Bias-cons, the consistency of the frequency bias between the different random seeded runs; (b) Bias-agg, the bias of the aggregate distribution (of a specific attribution method) over 10 fine-tuned models trained with different random seeds; (c) Bias-attr, the bias that is specific to the attribution method, compared to other methods. In principle, the framework is applicable to any attribution method and top-$k$ value.
@@
We investigate to what extent biases are consistent across different random seeds. Since we are interested in the distribution of top-$k$ selections over input tokens, we measure the mean pairwise Jensen-Shannon distance (Bias-cons) between the normalised distributions yielded by different runs. For this, we consider the unique model pairs $(m^{\prime},m^{\prime\prime})$ formed from the set of different runs $\mathcal{M}$ and we average the JS distance over the number of combinations, i.e. $\binom{|\mathcal{M}|}{2}=\frac{|\mathcal{M}|(|\mathcal{M}|-1)}{2}$. JS ranges from 0 (least divergent/biased) to 1 (most divergent/biased). The greater the distance, the greater the impact of random initialisation on the explanations. Formally, for each model type $m\in\{\text{{BERT, ModernBERT}}\}$ and each attribution method $a\in\mathcal{A}$, let $P^{(m^{\prime})}_{m,a}$ or $P^{(m^{\prime\prime})}_{m,a}$ denote the bias distribution (over positions or lexical elements in the input) for a single run $m^{\prime}$ or $m^{\prime\prime}$ (i.e., a specific random seed):
@@
We investigate the extent to which attribution methods are biased in the light of a model-to-model comparison. For each model type $m$, we compute the Jensen-Shannon distance (JS) between each method $a$’s aggregate bias distribution $P_{m,a}$ and a baseline. $P_{m,a}$ is computed by summing token-level bias distributions across all runs $\mathcal{M}$ and normalising to form a probability distribution. The baseline consists of a normalised uniform distribution $U$ over the positions and lexical elements for the specific dataset, and implies the absence of bias. $U$ is valid in a controlled setting and might need adjustment based on the nature of the task. We use a different baseline for causal relation detection in §4. We refer to this distance metric as Bias-agg, the bias of the aggregated distributions. Aggregate also refers to the fact that this bias conflates both model bias and attribution method bias.
@@
If the model were the only source of bias, all attribution methods would exhibit this bias consistently for the same model, same random seed. This is not the case (attributions differ across methods), indicating that at least some attribution methods introduce bias independently from the model. We define attribution bias (Bias-attr) as the averaged pairwise JS distance between the target method $a$ and each other method $a^{\prime}$ sampled from $\mathcal{A}$.
@@
It is unknown which of the attribution methods in $\mathcal{A}$ are the most biased. Based on the principle of reproducibility in science (Popper, 2005), a finding that is confirmed through different means is more likely to be correct: can a majority consensus among methods be indicative of faithfulness? Are outlier explanations more likely to be wrong? To answer these open questions, we investigate whether Bias-attr and Bias-agg behave similarly. In other words, are attribution methods that differ mostly from the rest (high Bias-attr) or are most similar to the rest (low Bias-attr) respectively the most and least biased with respect to a baseline expressing absence of bias (high and low Bias-agg)?
Faithfulness metrics: a sanity check
@@
We assess faithfulness through two widely used metrics: sufficiency (suff) (Lei et al., 2016), which is obtained by removing the top-1 tokens (the rationale $r_{i}$) from the input $x_{i}$ and observing the change in probability towards the ground truth label; comprehensiveness (cmp) (Yu et al., 2019), which is the change in probability obtained by removing the context ($x_{i}\setminus r_{i}$) around the top-1 tokens from $x_{i}$. Formally, following DeYoung et al. (2020):
@@
Most of the time, expecting a model to do better merely based on its size is not enough. To speculate more precisely how our models may relate to bias, we take into account the following three considerations. (i) Ben Amor et al. (2024) directly relate a form of position bias to the type of positional embeddings in the model. Contrary to BERT, which learnt fixed positional embeddings, ModernBERT uses Rotary Positional Embeddings (RoPE): along with the larger model size, this may lead to a reduced position bias. (ii) Lexical bias may also be reduced by RoPE in ModernBERT, giving more weight to the relative position of a lexical element in the context of the sentence, hence affecting its meaning. (iii) The effect on lexical bias caused by the greater diversity of datasets used by ModernBERT in pre-training is not univocal: it may either mitigate or enforce it.